CC12:

What is the equation of a line that is perpendicular to 2x+y=−4 and passes through the point (2, −8) ?

CC12:

\[y-(-8)=1/2(x-2)\]

CC12:

\[y=1/2(x-2)-8\]

CC12:

Answer \[y=1/2x-9\]

Shadow:

\[2x + y = -4\] \[y = -2x - 4\] Slope: -2, Perpendicular -> 1/2 Good It seems your x and y went in the right spots

Shadow:

1/2 times -2 = -1 Though it becomes +8 on the left side

Shadow:

So -1 - 8 = -9, looks good

CC12:

Thank you shadow